OpenStudy (anonymous):

find fourier coefficients discrete time x[n]=cos((2pi/5)n)+3sin((4pi/3)n)

OpenStudy (unklerhaukus):

\[x[n]=\cos(\frac{2\pi}{5}n)+3\sin\left(\frac{4\pi}{3}n\right)\]

OpenStudy (unklerhaukus):

what are we to find?

OpenStudy (unklerhaukus):

\[a_n=\frac 1 l \int\limits_a^{a+2l}f(x)\cos\frac{n\pi l}{x}\text d x\]

OpenStudy (anonymous):

In discrete time?

OpenStudy (unklerhaukus):

i dont get that part of the question

OpenStudy (anonymous):

Ok, find the fourier coefficients in discrete time

OpenStudy (anonymous):

I have the period: T =15 and I think taht the best way is through Euler's formula ¿?

OpenStudy (unklerhaukus):

if this question is not done using the formula i wrote above i'm sorry but i dont think i can help